# Photon statistics of resonantly driven spectrally diffusive quantum emitters

### v1

@@ V Acknowledgments
The authors thank Clarisse Fournier for fruitful discussions. This work is supported by the French Agence Nationale de la Recherche (ANR) under reference ANR-21-CE47-0004-01 (E--SCAPE project).
@@
Under non-resonant excitation, the emitted light does not exhibit bunching in the absence of additional processes causing intensity fluctuations, such as blinking. However, if a narrowband filter is placed on the detection path, spectral diffusion brings the emitter in and out of resonance with the filter, yielding intensity fluctuations after the filter. This technique shares some similarities with resonant excitation, on account of a spectrally selective process converting spectral variations into intensity fluctuations, which can be then analyzed using photon counting techniques. It has been used for instance to study SD in self-assembled quantum dots Abbarchi12 . For these reasons, in this section we investigate the statistical properties of filtered spontaneous emission of a spectrally diffusive emitter subject to either previously described SD mechanism.
@@
Experimentally, a measurement of the bunching time with at least two different filter widths would allow to infer the continuous nature of the spectral diffusion. In reference Abbarchi12 , since only one narrow filter was used to infer the bunching time, the study cannot be conclusive in this respect – and therefore the observed bunching time cannot be matched with the spectral diffusion time.
@@
Fig. S4 shows the detuning dependence of the zero-time bunching $B(0)$, as calculated from Eq. 3, for various SD amplitudes ranging from weakly broadened emitters (such as in ref. Kuhlmann13 ) to strongly broadened emitters (such as in refs. Fournier23PRB ; Wolters13 ). The detuning dependence shifts from a squared dispersive shape to an inverse Gaussian shape corresponding to the limit of Eq. 5. When the SD is weak, the fluorescence intensity is first-order insensitive to spectral fluctuations for exactly resonant laser drive (i.e. at zero detuning). In turn, the sensitivity is maximal at half-linewidth detuning – an observation already made in prior work on weakly broadened QDs Kuhlmann13 . On the other hand, in the case of strongly broadened emitters, the sensitivity sizably increases in the regions where the count rate vanishes. In this case, it is therefore more practical to perform experiments at resonance.
@@
hBN B center Fournier23PRB 88 MHz 4 GHz 30 $\mu$s 250
hBN red-emitting SPE Spokoyny20 $\mu$eV 50 $\mu$eV 23-510 $\mu$s 290-6400
NV center in diamond Wolters13 13 MHz 367 GHz 4.6 $\mu$s 1160
Self-assembled QDs Kuhlmann13 0.93 $\mu$eV 1.7 $\mu$eV 100 $\mu$s 8000
CdSe nanocrystals Beyler13 20 $\mu$eV 200 $\mu$eV 1 ms 150
Perovskite nanocrystals Utzat19 2.4-3.4 $\mu$eV 2 meV 0.5 ms 2000
GaN QDs Gao19 160  MHz 1.7 meV 260 ns 2200
@@
In this section, we consider an emitter subject to two SD processes with different timescales. Such situation has been encountered in prior work Spokoyny20 ; Abbarchi12 . We performed simulations by including two contributions to the SD – a combination of continuous and discrete diffusion, and a combination of two discrete processes. Fig. S6 shows the simulation results for the bunching function $\tilde{B}(\tau)$, where two decays can be observed in both cases. In the first case, both an exponential and a sub-exponential decay can be observed, while in the second case the decay is biexponential. Although we did not perform a full theoretical analysis, we observe that $\tilde{B}(\tau)$ can be fitted by a sum of both contributions, using Eqs. 8-9 for the OU process, and exponential decays for GJR processes.

### v2

@@ V Acknowledgments
The authors thank Clarisse Fournier for fruitful discussions. This work is supported by the French Agence Nationale de la Recherche (ANR) under reference ANR-21-CE47-0004-01 (E-SCAPE project).
@@
Under non-resonant excitation, the emitted light does not exhibit bunching in the absence of additional processes causing intensity fluctuations, such as blinking. However, if a narrowband filter is placed on the detection path, spectral diffusion brings the emitter in and out of resonance with the filter, yielding intensity fluctuations after the filter. This technique shares some similarities with resonant excitation, on account of a spectrally selective process converting spectral variations into intensity fluctuations, which can be then analyzed using photon counting techniques. It has been used for instance to study SD in self-assembled quantum dots [8]. For these reasons, in this section we investigate the statistical properties of filtered spontaneous emission of a spectrally diffusive emitter subject to either previously described SD mechanism.
@@
Experimentally, a measurement of the bunching time with at least two different filter widths would allow to infer the continuous nature of the spectral diffusion. In reference [8], since only one narrow filter was used to infer the bunching time, the study cannot be conclusive in this respect – and therefore the observed bunching time cannot be matched with the spectral diffusion time.
@@
Fig. S4 shows the detuning dependence of the zero-time bunching $B(0)$, as calculated from Eq. 3, for various SD amplitudes ranging from weakly broadened emitters (such as in ref. [24]) to strongly broadened emitters (such as in refs. [10, 30]). The detuning dependence shifts from a squared dispersive shape to an inverse Gaussian shape corresponding to the limit of Eq. 5. When the SD is weak, the fluorescence intensity is first-order insensitive to spectral fluctuations for exactly resonant laser drive (i.e. at zero detuning). In turn, the sensitivity is maximal at half-linewidth detuning – an observation already made in prior work on weakly broadened QDs [24]. On the other hand, in the case of strongly broadened emitters, the sensitivity sizably increases in the regions where the count rate vanishes. In this case, it is therefore more practical to perform experiments at resonance.
@@
[…]
GaN QDs [32] 160  MHz 1.7 meV 260 ns 2200
@@
In this section, we consider an emitter subject to two SD processes with different timescales. Such situation has been encountered in prior work [29, 8]. We performed simulations by including two contributions to the SD – a combination of continuous and discrete diffusion, and a combination of two discrete processes. Fig. S6 shows the simulation results for the bunching function $\tilde{B}(\tau)$, where two decays can be observed in both cases. In the first case, both an exponential and a sub-exponential decay can be observed, while in the second case the decay is biexponential. Although we did not perform a full theoretical analysis, we observe that $\tilde{B}(\tau)$ can be fitted by a sum of both contributions, using Eqs. 8-9 for the OU process, and exponential decays for GJR processes.